What is the yield to maturity of a five​-year, $ 5,000 bond with a 4.7​% coupon rate and semiannual coupons if this bond is currently trading for a price of $ 4, 787​?

A. 2.85​%

B. 7.97​%

C. 5.69​%

D. 6.83​%
Answer #1

C. 5.69​%

Yield-to-maturity = =rate(nper,pmt,-pv,fv)*2
= 5.69%
Where,
nper = 5*2 = 10
pmt = 5000*4.7%*6/12 = $     117.50
pv = $ 4,787.00
fv = $ 5,000.00